OpenStudy (anonymous):

determine the type of numbers of solution of 6x^2-7x-4=0

OpenStudy (anonymous):

imaginary numbers

OpenStudy (anonymous):

as said by dheeraj :)

OpenStudy (anonymous):

discriminant here is b^2 - 4ac = 49 - (4*4*6) = 49 - 96.. this is a negative number. So roots are imaginary

OpenStudy (anonymous):

thanks